You invested $1,000 two years ago, and the value of your investment has increased to $1,204. What is your compounded annual rate of return over this period?

a. 6.82%

b. 7.88%

c. 9.73%

d. 10.20%

e. 11.76%
Answer #1

c. 9.73%

A = P*(1+i)^n Where,
$    1,204 = 1000*(1+i)^2 A = $    1,204
1.204 = (1+i)^2 P = $    1,000
1.204 ^ (1/2) = 1+i i = ?
1.097269 = 1+i n = 2
0.097269 = i
So,
Annual rate of return is 9.73%